# To Shuffle or not to Shuffle: Auditing DP-SGD with Shuffling

### v2

@@ 6.3.2 Varying batch size $B$ and theoretical $\varepsilon$
Figure 6: Auditing DP-SGD (Shuffle) at various batch sizes $B$ and privacy levels $\varepsilon$.
@@
Next, we evaluate the impact of the batch size and theoretical (Poisson) $\varepsilon$ on the empirical privacy leakage observed from DP-SGD (Shuffle) under Partially-Informed. Due to computational constraints, rather than a CNN on CIFAR-10, we audit a LeNet model on FMNIST and an MLP model on P100 over $10^{6}$ observations. In Figure 6, we plot the empirical leakage estimates $\varepsilon_{emp}$ observed from auditing DP-SGD (Shuffle) at various batch sizes and privacy levels $\varepsilon$.
@@
With $B=100$, the empirical privacy leakage is significantly larger than the theoretical guarantees across both datasets and privacy levels. Specifically, for $B=100$ and theoretical $\varepsilon=1.0,2.0,4.0$, the models trained on FMNIST and P100 had empirical privacy guarantees of $\varepsilon_{emp}=3.46,4.66,5.78$ and $3.42,4.48,5.55$, respectively.
@@
However, with a larger batch $B=1000$, we only detect small gaps in some settings. Under Partially-Informed, the gradients of all samples remain unchanged except for the target record and final record in each batch. Thus, this likely introduces a “bias” term when evaluating the output from each step, i.e., $o^{i}_{t}\leftarrow\langle\tilde{g},\hat{g}\rangle$. Although Nasr et al. [43] show that this bias does not affect audits of DP-SGD (Poisson), they simply threshold the output from each step. Our auditing procedure is more complicated as we compute a likelihood ratio function across multiple steps, which we believe is more susceptible to the bias term thus resulting in weaker audits than expected from BGM. However, in several settings, the empirical privacy leakage is close to or already exceeds the theoretical Poisson guarantee, suggesting that given enough observations ($\approx 10^{9}$), we can potentially detect larger. Due to computational constraints, we leave this as part of future work.
6.3.3 Varying adversarial capabilities
@@
Our experiments thus far have shown that auditing DP-SGD (Shuffle) under Partially-Informed reveals large gaps between the empirical privacy leakage observed and the theoretical guarantees promised by the Poisson sub-sampling analysis. However, Partially-Informed is a stronger threat model than typically used in prior work [43]. Consequently, we set out to evaluate the impact of adversarial capabilities on the empirical privacy leakage observed.
@@
In Figure 7, we report the empirical privacy leakage estimates $\varepsilon_{emp}$ when auditing DP-SGD (Shuffle) under the three different threat models. We fix the batch size to $B=100$ and use $10^{6}$ observations. Despite the “bias” from other training samples discussed above, the empirical privacy leakage under Partially-Informed is almost the same as that from the idealized Worst-Case threat model (which corresponds to auditing BGM) at small batch sizes. Specifically, on FMNIST, at $\varepsilon=1.0,2.0,4.0$, auditing under Partially-Informed and Worst-Case yield privacy leakage estimates of $\varepsilon_{emp}=3.55,4.66,5.78$ and $\varepsilon_{emp}=4.06,5.18,6.43$, respectively.
@@
However, estimates under Natural are significantly weaker than the other two threat models. In all settings, $\varepsilon_{emp}$ under Natural is actually smaller than the theoretical $\varepsilon$ guaranteed by Poisson sub-sampling, even though shuffling was used. This is because the negative canary gradients inserted by the Partially-Informed adversary are crucial in “identifying” the batch containing the target (or zero-out) record in each epoch. Specifically, in the Partially-Informed threat model, without any noise, the adversary aims to distinguish between $(-1,...,+1,...,-1)$ and $(-1,...,0,...,-1)$, which has a Hamming distance of 2 with high probability (assuming the ‘+1’ and ‘0’ appear in different batches). Whereas under Natural, they have to distinguish between $(0,...,+1,...,0)$ and $(0,...,0)$, which only has a Hamming distance of 1. Informally, the distributions of outputs observed under Natural are less distinguishable than under Partially-Informed, thus resulting in lower empirical privacy leakage estimates.
6.4 Main Takeaways
Our experiments show that the empirical privacy leakage from shuffling is substantially larger than the theoretical guarantees stemming from the Poisson sub-sampling analysis. This gap occurs in a wide range of parameter settings studied and, crucially, in those used to train SOTA as well [19, 40]. For instance, De et al. [19] reportedly use shuffling to fine-tune a NF-ResNet-50 model on the Places-365 dataset at a theoretical $\varepsilon=7.53$, but our audits show that the actual privacy leakage is almost double ($\varepsilon_{emp}=13.59$).
We also show that the gap is also observable when auditing real models trained using the DP-SGD (Shuffle) algorithm. Specifically, when auditing a LeNet-5 classifier trained on the FMNIST dataset at theoretical $\varepsilon=1.0$ under the slightly weaker Partially-Informed threat model, we observe an empirical $\varepsilon_{emp}=3.46$, almost 3.5×3.5$\times$3.5 × the theoretical guarantee. However, under the weakest threat model considered, i.e., Natural, the $\varepsilon_{emp}$ is often below the expected theoretical guarantees $\varepsilon$, which suggests that strong threat models are necessary to audit shuffling mechanisms effectively.
@@ 7 Debugging Shuffle Implementations
Implementing DP algorithms correctly is known to be challenging [10, 34], and DP violations have been found in the wild [54, 4]. Specific to shuffling, Ponomareva et al. [49] report that for computational reasons, in many cases, datasets may not even be shuffled fully but only within a small buffer. Therefore, bugs or variations to the shuffling procedure itself may be present in DP-SGD implementations used in the wild, and, naturally, this would substantially affect their privacy leakage.
Based on a search of public GitHub code repositories (see Section 7.2), we identify cases where the dataset is first batched, and the batches are then shuffled instead of shuffling the samples before batching. As visualized in Appendix D.2, this makes the batch with the target record substantially more identifiable in cases where the samples surrounding it are different from samples “far away” from it (in this setting, batching is done locally). While these variations to the shuffling procedure intuitively affect the privacy analysis, in this section, we verify whether our auditing procedure can identify them and determine the scale of their impact on privacy leakage.
@@ 7.1 Partial Shuffling
1:Dataset, $D$. Number of observations, $N$. Target record, $(\hat{x},\hat{y})$. Zero-out record, $(x_{\bot},y_{\bot})$. Significance level, $\alpha$. Privacy parameter, $\delta$.
@@
7:         $[\mathbf{o}_{1}|...|\mathbf{o}_{T}]\leftarrow\Theta[i]$
8:         $[\mathbf{o^{\prime}}_{1}|...|\mathbf{o^{\prime}}_{T}]\leftarrow\Theta^{\prime}[i]$
@@
We first consider the case where datasets are only shuffled within a small buffer of $K$ samples, as reported in [49], and report the procedure used to debug this variation in Algorithm 2. The dataset is shuffled in buffers of size $K$, i.e., the first $K$ samples are shuffled and batched, followed by the next $K$ samples, etc. (See Appendix D.1 for a visualization of a small dataset that is partially shuffled.)
@@
Figure 8: Auditing DP-SGD (Shuffle) when the dataset is only shuffled within a buffer of $K$ samples. When the buffer $K=|D|=10{,}000$, the dataset is fully shuffled.
Results. In Figure 8, we report the empirical privacy leakage estimates for various buffer sizes $K$. We only audit the MLP model trained on P100 due to computational constraints. We experiment with small $\varepsilon=0.1,0.2,0.4$ as the violations are much more significant in this regime. Specifically, at $\varepsilon=0.1,0.2,0.4$, the empirical privacy leakages for full dataset shuffling $K=10{,}000$ and partial shuffling at $K=1{,}000$ are $\varepsilon_{emp}=0.09,0.31,1.80$ and $\varepsilon_{emp}=0.29,1.12,2.51$, respectively. This shows that even partial shuffling can result in substantially larger empirical privacy leakages.
However, as $\varepsilon$ increases, the gap between full shuffling and no shuffling ($K=100$) reduces. For instance, at $\varepsilon=0.1,0.2,0.4$, the $\varepsilon_{emp}$ values from no shuffling are 10.4$\times$, 6.16$\times$, and 1.72$\times$ that of the $\varepsilon_{emp}$ values from full shuffling, respectively. This indicates that for larger values of $\varepsilon$, the impact of not shuffling or partial shuffling steadily decreases, which in turn suggests that shuffling may not always result in strong privacy amplification.
@@ 7.2 Batch-then-Shuffle
Under the Natural and Partially-Informed threat models, batch-then-shuffle is equivalent to shuffle-then-batch. This is because we expect the dot product of the privatized gradient and gradient of almost all other samples, which cannot be modified by the adversary, to be near 0 under both threat models. Thus, this bug can only be detected under the Worst-Case threat model.
We then consider the Worst-Case threat model, but we let the adversary insert the gradients of the first $B$ samples as the canary gradient $\hat{g}$ and the remaining samples as the negative canary gradient $-\hat{g}$. By doing so, the outputs from the DP-SGD algorithm are expected to be $(B,-B,...,-B)+\mathcal{N}(0,\sigma^{2}\mathbb{I})$ and $(B-1,-B,...,-B)+\mathcal{N}(0,\sigma^{2}\mathbb{I})$ for $D$ and $D^{\prime}$, respectively. Note that ‘$B$’ and ‘$B-1$’ can appear uniformly in any batch. (See Appendix D.2 for a visualization of a small dataset that is batched first and then shuffled.)
Figure 9: Auditing DP-SGD (Shuffle) when the dataset is batched before shuffling is performed.
@@
7.3 Take-aways
The analysis presented in this section shows that our auditing procedure is highly extensible as it can be used to audit common variations of the shuffling procedure, more precisely, “partial shuffling” and “batch-then-shuffle.”
We show that these variations remain easily detectable by our auditing method, which estimates $\varepsilon_{emp}=0.29,1.12,2.51$ and $\varepsilon_{emp}=1.00,1.86,3.06$ for the “partial shuffling” and “batch-then-shuffle” procedure at theoretical $\varepsilon=0.1,0.2,0.4$, respectively.
@@ 8 Related Work
DP-SGD Audits. Jayaraman et al. [33] present the first DP-SGD audit using black-box inference attacks on the final trained model but only achieve loose estimates of the privacy leakage. Jagielski et al. [32] use input canaries to improve tightness and use Clopper-Pearson intervals to calculate confidence intervals for the empirical estimates. Nasr et al. [44] are the first to achieve tight DP-SGD audits using a large number of training runs, a pathological dataset, and an active white-box adversary that can insert gradient canaries at each step. To reduce the number of training runs, Zanella-Béguelin et al. [61] use credible intervals instead of Clopper-Pearson. Although [43] later questions the validity of the credible intervals, it shows that tight empirical estimates are possible even with few training runs and natural (not adversarially crafted) datasets by auditing with $f$-DP. Recent work on DP-SGD audits has also focused on federated learning settings [41, 2] and further reducing the number of training runs [48] to a single one [2, 52, 42] and auditing under weak threat models [3, 11].
@@
Overall, prior work auditing DP-SGD implementations has focused on Poisson sub-sampling. Previous audits also typically threshold the losses [3, 32, 42, 43, 44, 48, 52] or gradients [42, 43, 44, 44, 48] directly. By contrast, we audit the shuffling setting and use likelihood ratio functions to audit DP-SGD.
@@
Shuffling and DP-SGD. To the best of our knowledge, there is limited prior work analyzing the privacy of DP-SGD with shuffling. Chua et al. [16] analyze a simplified version of DP-SGD, i.e., the Adaptive Batch Linear Query (ABLQ) mechanism. Although they do not present a theoretical upper bound for the ABLQ mechanism with shuffling, they consider pathological settings where ABLQ could leak more privacy when using shuffling instead of Poisson sub-sampling. While they initially only consider a single epoch [16], in follow-up work [17], they extend their lower bound to multiple epochs. However, neither work directly analyzes DP-SGD or evaluates the gap in real-world parameter settings.
@@
Our work takes a different approach, empirically estimating the privacy leakage from DP-SGD using DP auditing. By doing so, we derive empirical privacy leakage estimates for real-world models trained using DP-SGD (Shuffle). We also focus on the impact of real-world considerations like threat models and variations on the shuffling procedure, which Chua et al. [16] mention would require more careful analysis but leave to future work.
@@
Shuffling. Shuffling is also used in other areas of DP, e.g., local DP, where users first randomize their inputs before submitting them for data processing. Bittau et al. [8] present the “Encode-Shuffle-Analyze” (ESA) framework where users’ randomized samples are shuffled before being processed. Erlingsson et al. [26] and Cheu et al. [12] formally prove that shuffling users’ data improves privacy guarantees in the local DP setting. Balle et al. [6] then improve on and generalize results from [26]. Erlingsson et al. [25] later revisit and apply their privacy amplification results to the ESA framework. More recently, Feldman et al. [27, 28] and Wang et al. [55] present a nearly optimal analysis of shuffling in the local DP setting. Overall, the privacy analysis for ($\varepsilon,\delta$)-DP mechanisms typically used in central DP is far less understood than that of shuffling for local DP mechanisms, which typically guarantee pure DP (i.e., $\delta=0$).
9 Conclusion

### v3

@@ 6.2.2 Varying batch size $B$ and theoretical $\varepsilon$
Figure 6: Auditing DP-SGD (Shuffle) at various batch sizes ($B$) and privacy levels ($\varepsilon$).
@@
Next, we evaluate the impact of the batch size and theoretical (Poisson) $\varepsilon$ on the empirical privacy leakage observed from DP-SGD (Shuffle). Due to computational constraints, rather than a CNN on CIFAR-10, we audit a LeNet model on FMNIST and an MLP model on P100 over $10^{6}$ observations. In Figure 6, we plot the empirical leakage estimates $\varepsilon_{emp}$ observed from auditing DP-SGD (Shuffle) at various batch sizes and privacy levels $\varepsilon$.
@@
With $B=100$, the empirical privacy leakage is significantly larger than the theoretical guarantees across both datasets and privacy levels. Specifically, for $\varepsilon=1.0,2.0,4.0$, privacy leakage from the models trained on FMNIST and P100 amount to, respectively, $\varepsilon_{emp}=3.46,4.66,5.78$ and $3.42,4.48,5.55$.
@@
However, with a larger batch $B=1000$, we only detect small gaps in some settings. Under Partially-Informed, the gradients of all samples remain unchanged except for the target record and final record in each batch. Thus, this likely introduces a “bias” term when evaluating the output from each step, i.e., $o^{i}_{t}\leftarrow\langle\tilde{g},\hat{g}\rangle$. Although Nasr et al. [42] show that this bias does not affect audits of DP-SGD (Poisson), they simply threshold the output from each step. Our auditing procedure is more complicated as we compute a likelihood ratio function across multiple steps, which we believe is more susceptible to the bias term, resulting in weaker audits than expected from BGM. However, in several settings, the empirical privacy leakage is close to or already exceeds the theoretical Poisson guarantee, suggesting that given enough observations ($\approx 10^{9}$), we can potentially detect larger. Due to computational constraints, we leave this to future work.
6.3 Varying Adversarial Capabilities
@@
Our experiments thus far show that auditing DP-SGD (Shuffle) under Partially-Informed reveals large gaps between the empirical privacy leakage observed and the theoretical guarantees promised by the Poisson subsampling analysis. Next, we set out to evaluate the impact of different adversarial capabilities on the empirical privacy leakage.
@@
In Figure 7, we report the different values of $\varepsilon_{emp}$ when auditing DP-SGD (Shuffle) with the three different adversaries considered (see Section 5). We fix the batch size to $B=100$ and use $10^{6}$ observations. Despite the “bias” from other training samples discussed above, the empirical privacy leakage under Partially-Informed is almost the same as that from the idealized Worst-Case threat model (which corresponds to auditing BGM) at small batch sizes. Specifically, on FMNIST, at $\varepsilon=1.0,2.0,4.0$, auditing under Partially-Informed and Worst-Case yield privacy leakage estimates, respectively, of $\varepsilon_{emp}=3.55,4.66,5.78$ and $\varepsilon_{emp}=4.06,5.18,6.43$.
@@
However, estimates under Target-Canary are significantly weaker than the other two threat models. In all settings, $\varepsilon_{emp}$ under Target-Canary is actually smaller than the theoretical $\varepsilon$ guaranteed by Poisson subsampling, even though shuffling was used. This is because the negative canary gradients inserted by the Partially-Informed adversary are crucial in “identifying” the batch containing the target (or zero-out) record in each epoch. Specifically, in the Partially-Informed threat model, without any noise, the adversary aims to distinguish between (1,,+1,,1)(-1,...,+1,...,-1) and (1,,0,,1)(-1,...,0,...,-1), which has a Hamming distance of 2 with high probability (assuming the ‘+1’ and ‘0’ appear in different batches). Whereas under Target-Canary, they have to distinguish between (0,,+1,,0)(0,...,+1,...,0) and (0,,0)(0,...,0), which only has a Hamming distance of 1. Informally, the distributions of outputs observed under Target-Canary are less distinguishable than under Partially-Informed, thus resulting in lower empirical privacy leakage estimates.
6.4 Takeaways
Our experiments confirmed that the empirical privacy leakage observed from shuffling is much larger than the theoretial privacy leakage expected from the Poisson subsampling analysis when auditing models trained using the DP-SGD (Shuffle) algorithm. Specifically, for a LeNet-5 classifier trained on FMNIST at theoretical $\varepsilon=1.0$ in the Partially-Informed model, we find an empirical privacy leakage amounting to $\varepsilon_{emp}=3.46$, almost $3.5\times$ the theoretical guarantee.
We also show that audits using a weaker adversary, i.e., Target-Canary, the $\varepsilon_{emp}$ is well below the expected theoretical guarantees $\varepsilon$, which confirms that strong adversarial models are necessary to audit shuffling mechanisms effectively.
@@ 7 Debugging Shuffle Implementations
Implementing DP algorithms correctly is known to be challenging [10, 33], and DP violations have been found in the wild [53, 4]. Specific to shuffling, Ponomareva et al. [47] report that for computational reasons, in many cases, datasets may not even be shuffled fully but only within a small buffer. Therefore, bugs or variations to the shuffling procedure itself may be present in DP-SGD implementations, and, naturally, this would substantially affect their privacy leakage.
Searching public GitHub repositories (see Section 7.2), we identify cases where the dataset is first batched, and the batches are then shuffled, rather than shuffling the samples before batching. As visualized in Section 7.2, this makes the batch with the target record substantially more identifiable in cases where the samples surrounding it are different from samples “far away” from it (in this setting, batching is done locally). While these variations to the shuffling procedure intuitively affect the privacy analysis, in this section, we verify whether our auditing procedure can identify them and determine the scale of their impact on privacy leakage.
@@ 7.1 Partial Shuffling
[…]
@@
7:    $\mathbf{o}_{1}|...|\mathbf{o}_{T}]\leftarrow\Theta[i]$
8:    $\mathbf{o^{\prime}}_{1}|...|\mathbf{o^{\prime}}_{T}]\leftarrow\Theta^{\prime}[i]$
@@
Figure 8: Comparison between a small dataset with eight records ($x_{1}$$x_{7}$$x_{T}$) being fully shuffled vs. partially shuffled with a buffer of K=4K{=}4 and batch size B=2B{=}2. The target record $x_{T}$ is highlighted in red.
We first consider the case where datasets are only shuffled within a small buffer of $K$ samples, as reported in [47], and report the procedure used to debug this variation in Algorithm 3. The dataset is shuffled in buffers of size $K$, i.e., the first $K$ samples are shuffled and batched, followed by the next $K$ samples, etc. Figure 8 provides a visualization of a small one-dimensional dataset that is partially shuffled.
@@
We review the procedure used to debug this variation in Algorithm 3. We denote with $\text{DP-SGD}^{PI}_{K}$ the execution of DP-SGD with partial shuffle over a buffer of size $K$ and with the modifications made for the Partially-Informed adversary as reported in Algorithm 2.
Figure 9: Auditing DP-SGD (Shuffle) when the P100 dataset is only shuffled within a buffer of $K$ samples. When the buffer K=|D|=10, [^0] K=|D|=10{,}000, the dataset is fully shuffled.
Results. In Figure 9, we report the empirical privacy leakage estimates for various buffer sizes $K$. We only audit the MLP model trained on P100 due to computational constraints. We experiment with small $\varepsilon=0.1,0.2,0.4$ as the violations are much more significant in this regime. Specifically, at $\varepsilon=0.1,0.2,0.4$, the empirical privacy leakages for full dataset shuffling K=10, [^0] K=10{,}000 and partial shuffling at K=1, [^0] K=1{,}000 are $\varepsilon_{emp}=0.09,0.31,1.80$ and $\varepsilon_{emp}=0.29,1.12,2.51$, respectively. This shows that even partial shuffling can result in substantially larger empirical privacy leakages.
@@ 7.2 Batch-then-Shuffle
Under the Target-Canary and Partially-Informed threat models, batch-then-shuffle is equivalent to shuffle-then-batch. This is because we expect the dot product of the privatized gradient and gradient of almost all other samples, which cannot be modified by the adversary, to be near 0 under both threat models. Thus, this bug can only be detected under the Worst-Case threat model. We then consider the Worst-Case threat model, but we let the adversary insert the gradients of the first $B$ samples as the canary gradient $\hat{g}$ and the remaining samples as the negative canary gradient $\hat{g}$. By doing so, the outputs from the DP-SGD algorithm are expected to be $(B,-B,...,-B)+\mathcal{N}(0,\sigma^{2}\mathbb{I})$ and $(B-1,-B,...,-B)+\mathcal{N}(0,\sigma^{2}\mathbb{I})$ for $D$ and $D^{\prime}$, respectively. Note that ‘$B$’ and ‘B1B-1’ can appear uniformly in any batch. Please refer to Figure 10 for a visualization of a small one-dimensional dataset that is batched first and then shuffled.
@@
7.3 Takeaways
Our experiments show that our auditing procedure is highly extensible as it can be used to audit common variations of the shuffling procedure, more precisely, “partial shuffling” and “batch-then-shuffle.” These variations remain easily detectable by our auditing method, which estimates $\varepsilon_{emp}=0.29,1.12,2.51$ and $\varepsilon_{emp}=1.00,1.86,3.06$ for the “partial shuffling” and “batch-then-shuffle” procedure at theoretical $\varepsilon=0.1,0.2,0.4$, respectively.
@@ 8 Related Work
DP-SGD Audits. Jayaraman et al. [32] present the first DP-SGD audit using black-box inference attacks on the trained model, but only achieve loose estimates of the privacy leakage. Jagielski et al. [31] use input canaries to improve tightness and use Clopper-Pearson intervals to calculate confidence intervals for the empirical estimates. Nasr et al. [43] are the first to achieve tight DP-SGD audits using a large number of training runs, a pathological dataset, and an active white-box adversary that can insert gradient canaries at each step. To reduce the number of training runs, Zanella-Béguelin et al. [59] use credible intervals instead of Clopper-Pearson. Although [42] later questions the validity of the credible intervals, it shows that tight empirical estimates are possible even with few training runs and natural (not adversarially crafted) datasets by auditing with $f$-DP. Recent work on DP-SGD audits has also focused on federated learning settings [40, 2] and further reducing the number of training runs [46] to a single one [2, 50, 41] and auditing under weak threat models [3, 11].
@@
Overall, prior work auditing DP-SGD implementations has focused on Poisson subsampling. Previous audits also typically threshold the losses [3, 31, 41, 42, 43, 46, 50] or gradients [41, 42, 43, 43, 46] directly. By contrast, we audit the shuffling setting and use likelihood ratio functions to audit DP-SGD.
Figure 11: Auditing DP-SGD (Shuffle) when the P100 dataset is batched before shuffling is performed.
@@
Shuffling and DP-SGD. To the best of our knowledge, there is limited prior work analyzing the privacy of DP-SGD with shuffling. Chua et al. [16] analyze a simplified version of DP-SGD, i.e., the Adaptive Batch Linear Query (ABLQ) mechanism. Although they do not present a theoretical upper bound for the ABLQ mechanism with shuffling, they consider pathological settings where ABLQ could leak more privacy when using shuffling instead of Poisson subsampling. While they initially only consider a single epoch [16], in follow-up work [17], they extend their lower bound to multiple epochs. However, neither work directly analyzes DP-SGD or evaluates the gap in real-world parameter settings.
@@
Our work takes a different approach, empirically estimating the privacy leakage from DP-SGD using DP auditing. By doing so, we derive empirical privacy leakage estimates for real-world models trained using DP-SGD (Shuffle). We also focus on the impact of real-world considerations like threat models and variations on the shuffling procedure. Overall, we believe that our work is orthogonal to [16] as our focus is on empirical privacy estimation, whereas theirs is on theoretical privacy analysis.
@@
Shuffling. Shuffling is also used in local DP, where users randomize their inputs before sharing them for data processing. Bittau et al. [8] introduce the Encode-Shuffle-Analyze (ESA) framework, where users’ randomized samples are shuffled before being processed. Cheu et al. [12] and Erlingsson et al. [26] prove that shuffling users’ data improves privacy guarantees in the local DP setting. In follow-up work, Erlingsson et al. [25] apply their privacy amplification results to the ESA framework, while Balle et al. [6] improve on and generalize results from [26]. Finally, Feldman et al. [27, 28] and Wang et al. [54] present a nearly optimal analysis of shuffling in the local DP setting. Overall, the privacy analysis for ($\varepsilon,\delta$)-DP mechanisms typically used in central DP is far less understood than that of shuffling for local DP mechanisms, which typically guarantee pure DP (i.e., $\delta=0$).
9 Discussion and Conclusion